PLEASE HELP ME!
1. The perimeter of a rectangle is 68 meters. The length is 2 meters less than triple the width. What are the dimensions of the rectangle?
(a) Define the variable expressions.

(b) Write the equation.
(c) Solve the equation and determine the values ased for.

Answer by smooth criminal(1)

data
perimetre- 68 metres
Let breadth be b
length will be 3b-2
2(3b-2+b)= 68
3b-2+b= 68/2
4b-2=34
4b=34+2=36
b=36/4
breadth is 9
length is 3(9)-2= 25